A and B which are separated by 920 Km are directly connected using some type of physical link. A is sending packets to B with a packet size of 8000 bits at a bit rate of 2 Mbps. If the first bit transmitted by A has just reached B when the last bit is being transmitted by A, calculate the velocity of propagation in the medium in two different ways:
(i) Using the concept of delay-bandwidth product
(ii) Using the concept of bit width
Answer #1

1 - Using the concept of delay-bandwidth product

distance = 920 km

Propagation time = the time taken by a bit to travel from source to destination

Round Trip Time = 2 * Propogation time.

BandWidth Product is the the amount of data that can be in transit in the network. which is

BandWidthProduct = total_available_bandwidth * roundTripTime

In this case Band WidthProduct = 8000 bits because the first bit transmitted by A has just reacched B when

the last bit is being transmitted by A.

=>  8000 = 2 * 10^{6} * (Round Trip Time) => 8000 = 2 * 10^{6} * 2 * \frac{distance}{velocity of propagation}

=>velocity of propagation = \frac{2 * 10^{6} * 2 * 920 * 10^{3}}{8000} = 0.46 * 10^{9} m/s

2 - Using the concept of bitwidth

Propogation delay = \frac{distance}{velocity of propogation}

In this case as mentioned the first bit transmitted by A has just reached B when the last bit is being transmitted by A

So the transmission time at A is equal to propogation delay

=> propogation delay = \frac{8000}{2 * 10^{6}} = 4 * 10^{-3} secs = 4 milli secs

=> 4 * 10^{-3} = \frac{920 * 10 ^{3}}{ velocity}

=> velocty = \frac{920 * 10 ^{3}}{4 * 10^{-3}} => 230 * 10 ^{6} m/sec => 2.30 * 10 ^{8}

that's it !

/* if any queries please comment */

/* thank you */